OpenStudy (anonymous):

A town has 1800 people initially. In each of the cases below, find the formula for the population of the town, P, in terms of the number of years, t.

OpenStudy (anonymous):

The town grows by 50 people a year. Answer: P(t)= The town shrinks by 80 people a year. Answer: P(t)=

OpenStudy (anonymous):

I know that when the town grows at an annual rate of 3% a year the answer is P(t)= 1800(1.03)^t

OpenStudy (anonymous):

And I also know that when the town shrinks at an annual rate of 5 percent a year the answer is P(t)= 1800(0.95)^t

OpenStudy (anonymous):

@misty1212 Any chance you can help?

OpenStudy (calculusxy):

Would the town growing by 50 people a year be P(t) = 1800 + 50t because it's 50 people/year. And then shrinking would just be P(t) = 1800 - 80t because when it's shrinking it's declining/subtracting and it's 80 people/year so 80 times "t".

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is what @calculusxy wrote, not exponential because they tell you 50 Per Year

OpenStudy (anonymous):

Yep that's correct! Thanks @calculusxy For some reason when I entered in my answer as 1800+(50)^t it was wrong but 1800+50t was correct!

OpenStudy (calculusxy):

You're welcome :)